OpenStudy (riley):

A Paramecium swims at 1mm/sec. In an hour it swims about 10^4 body lengths. a) How long is a Paramecium? b) How fast in miles per hour is the paramecium swimming?

OpenStudy (trancenova):

a) 1hour = 3600 seconds, v (velocity)= d (distance)/t (time), d=v*t. d=1mm/sec * 3600. d=3600 mm So for body length you need 10^4/3600 mm = 2.78mm, so your paramecium is about 2.78mm.